Dear sir, Help me solving the below problem. The cost of fuel for running a train is proportional to the square of the speed in kmph, & costs Rs75 per hour at 17kmph. What is the most economical speed, if the fixed charges are Rs400 per hour Regards Girish

Dear sir,


Help me solving the below problem.


The cost of fuel for running a train is proportional to the square of the speed in kmph, & costs Rs75 per hour at 17kmph. What is the most economical speed, if the fixed charges are Rs400 per hour


 


Regards


Girish

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear girish

first we should calculate the cost function

C=kv2

put the condition at v=17  ,C= 75

 so k =75/(17)2

so C = 75(v/17)2

now let train has to travel L km

 so time taken by train t =L/v hr

so total cost CT = [75(v/17)2]L/v   +  400 * L/v

                     = 25L [ 3V /172 + 8/V ]

for minimum cost  dCT/dv =0

    this give V = 34√2/√3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin